OpenStudy (anonymous):

Solve the equation for x. Round your answer to two decimal places, and do not include "x =" in your answer. 3*ln x = 7.6

OpenStudy (anonymous):

0.93

OpenStudy (anonymous):

Sure ? I got ln x = 7.6/3 x = e^(7.6/3) x= 12.5954

OpenStudy (anonymous):

wiliam was right. its 12.60